# A Unified Approach to Submodular Maximization Under Noise

### v1

@@ Abstract
We consider the problem of maximizing a submodular function with access to a noisy value oracle for the function instead of an exact value oracle. Similar to prior work [16, 13], we assume that the noisy oracle is persistent in that multiple calls to the oracle for a specific set always return the same value. In this model, Hassidim and Singer [13] design a (11/e)(1-1/e)-approximation algorithm for monotone submodular maximization subject to a cardinality constraint and Huang et al. [16] design a (11/e)/2(1-1/e)/2-approximation algorithm for monotone submodular maximization subject to any arbitrary matroid constraint. In this paper, we design a meta-algorithm that allows us to take any “robust” algorithm for exact submodular maximization as a black box and transform it into an algorithm for the noisy setting while retaining the approximation guarantee. By using the meta-algorithm with the measured continuous greedy algorithm, we obtain a (11/e)(1-1/e)-approximation (resp. $1/e$-approximation) for monotone (resp. non-monotone) submodular maximization subject to a matroid constraint under noise. Furthermore, by using the meta-algorithm with the double greedy algorithm, we obtain a $1/2$-approximation for unconstrained (non-monotone) submodular maximization under noise.
@@ 1 Introduction
The standard model in submodular maximization assumes that we have value oracle access to the submodular function $f$ where, given a set $S$, we can retrieve the exact value of $f(S)$. However, in many settings it is not realistic to assume the existence of an exact value oracle. For example, in machine learning, we can never evaluate the true loss function of a model as we do not have access to the true distribution of the population. At best, we may have access to a noisy version of the loss function. In this paper, we study a model for noisy submodular maximization introduced by Hassidim and Singer [13] where querying a set $S$ returns only an unbiased estimate of $f(S)$. A notable feature of this model is persistent noise, where querying the value of a set twice returns the same value. Thus, simply querying a submodular function multiple times at the same input cannot be used to denoise the function. Nonetheless, one can still ask whether existing algorithms work in this setting. However, Hassidim and Singer [13] show that this is probably unlikely. For example, they show that the natural greedy algorithm is too sensitive to noise and only obtains an $o(1)$-approximation despite being the optimal algorithm when there is no noise. Thus, new algorithms are needed to deal with noise.
For the problem of maximizing a noisy monotone submodular function subject to a cardinality constraint, Hassidim and Singer [13] designed an algorithm that achieves a tight (11/e)(1-1/e)-approximation. At a high-level, their algorithm computes a smooth surrogate function which essentially averages a few queries together to obtain an estimate that may have less noise but potentially be biased. They then apply the greedy algorithm on this surrogate function to obtain a (11/e)(1-1/e)-approximation. In a follow-up work, Huang et al. [16] proved a (11/e)/2(1-1/e)/2-approximation ratio for noisy monotone submodular maximization under an arbitrary matroid. This approximation ratio can be improved to 11/e1-1/e if the matroid satisfies a “strongly base-orderable” property. While Huang et al. [16] also make use of the smoothing technique, their algorithm is based on a local search algorithm [10].
@@
While prior work has established that submodular maximization is feasible in the noisy setting, a notable downside is that the algorithms and analyses are designed for their specific problems. For example, Hassidim and Singer [13] only obtain results for monotone submodular maximization subject to a cardinality constraint and, while Huang et al. [16] extend to arbitrary matroids, their analyses have a factor of $2$ gap with the optimal result, despite an algorithm which is provably optimal in the non-noisy setting. Our main question is whether such specific analyses are necessary. In particular, we want to answer the following question:
@@
Our main contribution in this paper is an affirmative answer to the above question. In particular, we show that any algorithm for exact submodular optimization which is sufficiently “robust” can be translated to an algorithm for noisy submodular optimization with only an $o(1)$ loss in the approximation ratio. For many settings, this obviates the need for designing new algorithms and instead relies only on checking whether existing algorithms are sufficiently robust. Our technique builds on but differs from [13]: we propose the use of a random surrogate function, which circumvents the limitation of a deterministic surrogate function in [13].
For our first application, we instantiate our framework with the continuous greedy algorithm [4, 9]. Specifically, we use the measured continuous greedy algorithm of Feldman et al. [9] which gives a (11/e)(1-1/e)-approximation for monotone submodular maximization and a $1/e$-approximation for non-monotone submodular maximization, both subject to matroid constraints. As our framework inherits the approximation guarantees of these algorithms, we achieve the same approximation ratios for the noisy setting, with only $o(1)$ loss in the approximation ratio. Our second application is to instantiate our framework with the double greedy algorithm [3] to obtain a $1/2$-approximation for noisy unconstrained submodular maximization, which is tight even without noise. To our knowledge, we provide the first set of results for noisy submodular maximization for non-monotone functions, in both unconstrained and matroid constrained settings. We summarize our results in Table 1.
@@
Another work that is closely related to ours is Hassidim and Singer [14], which also studies the setting where the noise is persistent. In particular, they considered intersection of $P$ matroids and show a $1/(1+P)$-approximation. There have also been works that look at the noisy submodular setting but without persistent noise: for example, Singla et al. [20] consider a model where one only has preference information, and Chen et al. [5] assume that one has access to noisy marginal estimates instead of a value oracle.
Another related line of work is optimizing approximately submodular functions, where the noise is not unbiased and can be adversarial. Horel and Singer [15] show that even there is only $1/\sqrt{n}$ noise, the problem becomes intractable and prove exponential lower bounds on the query complexity. On the other hand, they show that constant approximation algorithms are achievable when the “curvature” is bounded or when the noise is small in terms of the rank of the matroid. Zheng et al. [25] study maximizing approximately $k$-submodular functions. Chierichetti et al. [6] study functions that satisfy submodularity $\varepsilon$-approximately and convert them to exactly submodular functions that are $\varepsilon n^{2}$-close to the original functions, establishing a reduction to the exact submodular setting. A difference with our work is that our noisy value function is far from being $\varepsilon$-approximately submodular. A key contribution of our work is to convert our noisy value function to a surrogate function that is close to the underlying submodular function, which then enables a reduction similar to [6].
A less related line of work is on online submodular maximization [22, 18, 21, 12, 17]. Here, one is given a sequence of submodular functions $f_{1},f_{2},\ldots$ and before seeing $f_{t}$, a player must decide on a set to play. The goal is to minimize $\alpha$-regret” which is the difference between the value obtained by the player and an $\alpha$-approximation of the optimum. Although the problem setting is different, these works also tend to show that many known algorithms are “robust” and develop a way to reuse these algorithms in the online setting. Such an approach is similar to ours as we show that a number of algorithms are robust and reuse them for the noisy setting.
@@ 2 Preliminaries
A noisy value oracle for $f$ is denoted by $\tilde{f}:2^{N}\to{\mathbb{R}}_{\geq 0}$. Following previous works [13, 16], we consider a multiplicative noisy value oracle defined by $\tilde{f}(S)=\xi_{S}f(S)$ for all sets $S\subseteq N$, where $\xi_{S}$ is a non-negative random variable distributed according to some distribution $\mathcal{D}$. We call $\xi_{S}$ the noise multiplier for set $S$. Following [13, 16], we assume that the noisy value oracle satisfies the following three properties:
@@
Definition 2.1 (see, e.g., Wainwright [24]).
@@
Sub-exponential distributions are a large class of distributions, including bounded, Gaussian, and exponential distributions. For example, a random variable $\xi$ bounded in $[0,B]$ is sub-exponential with parameters $(\nu=B,\alpha=0)$. Moreover, sub-exponential distributions include the generalized exponential tail distributions that have been considered by previous work on monotone submodular maximization under noise [13, 16]. We assume that the parameters $(\nu,\alpha)$ of the sub-exponential distribution $\mathcal{D}$ are known while the distribution $\mathcal{D}$ itself is unknown.
@@ 3 A Unified Approach to Noisy Submodular Maximization
Our reduction uses an idea proposed by previous work on noisy submodular maximization [13, 16]. This idea is to estimate the value $f(S)$ of a set $S$ using the noisy value of some surrogate function $F(S)=\frac{1}{|\mathcal{T}_{S}|}\sum_{T\in\mathcal{T}_{S}}f(T)$ where $\mathcal{T}_{S}$ is a collection of sets related to $S$. One example is to choose a small set $H$ and let $\mathcal{T}_{S}=\{S\cup H^{\prime}\text{ for $H^{\prime}\subseteq H$}\}$. In previous work, different surrogate functions were constructed for different settings. Here, we construct a single surrogate function for all the settings we consider (unconstrained non-monotone maximization, monotone and non-monotone maximization under matroid constraints); see Section 3.1 for details. Then, the reduction is to run the algorithm $\mathcal{A}$ to maximize the noisy surrogate function $\tilde{F}(S)=\frac{1}{|\mathcal{T}_{S}|}\sum_{T\in\mathcal{T}_{S}}\tilde{f}(T)\approx F(S)$.
The main technical challenge here is to ensure that the surrogate function $F(S)$ approximates the true function $f(S)$ well. As shown by [13, 16], a deterministic surrogate function does not always guarantee a good approximation to the true function. So, they need to use arguments that are specific to their algorithms to show that the surrogate functions work well. A key insight in our work is the use of a random surrogate function. We show that such a random surrogate function can approximate the true function in all settings with matroid constraints (see Section 3.3).
@@ 3.1 Surrogate Function
Let $H\subseteq N$ be a subset of size $|H|=h$ where $h$ is a small integer. We call $H$ a smoothing set. Let $t<h$ be another integer. Let $H[t]=\{H^{\prime}\subseteq H:|H^{\prime}|=t\}$ be all the subsets of $H$ of size $t$; there are $\binom{h}{t}$ such subsets. We use $H^{\prime}\sim H[t]$ to denote the random sample of a subset $H^{\prime}\subseteq H$ of size $t$ uniformly at random. Define a surrogate function $F^{H,t}$ as follows:
@@ 3.2 A Meta-Algorithm for Noisy Submodular Maximization
Lemma 3.6 (Donald and Tobey [7]).
@@ 4.1 Matroid Constraints
We first consider maximizing monotone and non-monotone submodular functions under matroid constratins. We prove that the measured continuous greedy algorithm of Feldman et al. [9] is robust. Here, we recall the algorithm (with exact value oracle). First, define the multilinear extension of a submodular function $f$ as
@@
where the $\operatorname*{arg\,max}$ is taken over the matroid polytope $\mathcal{P}$. Then we update $x_{i}(t+\delta)=x_{i}(t)+\delta(1-x_{i}(t))y_{i}^{*}(t)$. Finally, we use pipage rounding [4] (which is an obvlious rounding scheme) to convert the fractional solution $x(1)$ to a discrete set $S\in\mathcal{I}$. The pipage rounding technique guarantees that $\mathbb{E}[f(S)]\geq F(x(1))$. In addition, we have that $F(x(1))/f(\operatorname{OPT})\geq 1-1/e-O(n^{3}\delta)$ when $f$ is monotone and $F(x(1))/f(\operatorname{OPT})\geq 1/e-O(n^{3}\delta)$ when $f$ is non-monotone.
@@
For submodular function maximization under matroid constraint, the measured continuous greedy algorithm [9] obtains a (11/e)(1-1/e)-approximation when $f$ is monotone and a $(1/e)$-approximation when $f$ is non-monotone. Moreover, the algorithm is $O(n\varepsilon)$-robust.
@@
We apply Theorem 3.4 to the problem of maximizing monotone submodular functions with matroid constraints under a noisy value oracle. Let the algorithm $\mathcal{A}$ in Algorithm 1 be the measured continuous greedy algorithm [9] mentioned above. Fix parameter $\varepsilon\in(0,\frac{2\nu^{2}}{\alpha})$. Choose integer $m\geq\max\{2,8\nu^{2}\}\frac{n^{4}}{\varepsilon^{2}}\big(n+\log(4n)\big)=\tilde{O}(\frac{n^{5}}{\varepsilon^{2}})$, $t\geq\log_{2}(4m)=\Theta(\log(\frac{n}{\varepsilon}))$, and $h=t^{2}=\Theta(\log^{2}(\frac{n}{\varepsilon}))$. Then, we apply Theorem 3.4 and Lemma 4.1 with parameter $\varepsilon_{1}=\frac{\varepsilon f_{\max}}{n^{2}}$ to obtain:
@@
Fix $\varepsilon\in(0,\frac{2\nu^{2}}{\alpha})$. Suppose $n\geq\frac{1}{\varepsilon}$ and the matroid’s rank $r\geq\Omega(\frac{1}{\varepsilon}\log^{2}(\frac{n}{\varepsilon}))$. By letting the $\mathcal{A}$ in Algorithm 1 be the measured continuous greedy algorithm [9], we obtain a polynomial-time algorithm for maximizing monotone submodular functions under matroid constraints with noisy oracle satisfying $\mathbb{E}[f(\operatorname{ALG})]~\geq~\big(1-\frac{1}{e}-O(\varepsilon)\big)f(O^{*})$.
Previous work [16] gave a polynomial-time algorithm with $(1-1/e)/2-O(\varepsilon)$ approximation ratio for noisy monotone submodular maximization under matroid constraints, assuming $n\geq\varepsilon^{-4}$ and $r\geq\varepsilon^{-4/3}$. Our Theorem 4.2 improves [16] by increasing the approximation ratio to $1-1/e-O(\varepsilon)$ (which is the tight ratio even with the exact value oracle) as well as relaxing the condition on $n$ and $r$.
In the special case of cardinality constraints, [16]’s algorithm achieves $(1-1/e-O(\varepsilon))$ approximation with $\tilde{O}(r^{2}n^{3}/\varepsilon)$ query complexity (their Theorem 4.6). By letting $\mathcal{A}$ be the greedy algorithm with query complexity $Q(\mathcal{A})=O(rn)$, our algorithm achieves the same approximation ratio with query complexity $m\cdot Q(\mathcal{A})=\tilde{O}(n^{5}/\varepsilon^{2})\cdot O(rn)=\tilde{O}(rn^{6}/\varepsilon^{2})$. While our query complexity might be worse than [16], the advantage of our algorithm lies in its generality and better approximation ratio in the more general matroid constraints case.
@@
Fix $\varepsilon\in(0,\frac{2\nu^{2}}{\alpha})$. Suppose $n\geq\frac{1}{\varepsilon}$ and the matroid’s rank $r\geq\Omega(\frac{1}{\varepsilon}\log^{2}(\frac{n}{\varepsilon}))$. By letting the $\mathcal{A}$ in Algorithm 1 be the measured continuous greedy algorithm [9], we obtain a polynomial-time algorithm for maximizing non-monotone submodular functions under matroid constraints with noisy oracle satisfying $\mathbb{E}[f(\operatorname{ALG})]~\geq~\big(\frac{1}{e}-\frac{1}{\Theta(\log\frac{n}{\varepsilon})}-O(\varepsilon)\big)f(O^{*})$.
The $\frac{1}{e}\approx 0.367$ approximation above is not necessarily tight. Buchbinder and Feldman [2] design a $0.401$-approximation algorithm for non-monotone submodular maximization under matroid constraint with exact value oracle. Their algorithm is quite technical, so we leave as an open question whether their algorithm is robust against approximate value oracle. If their algorithm is robust, then it can be directly converted to an algorithm for noisy non-monotone submodular maximization under matroid constraint achieving 0.401o(1)0.401-o(1) approximation, using our Algorithm 1.
@@
While our main results (Theorems 3.4, 4.2 and 4.3) are stated in terms of the expected value $\mathbb{E}[f(\operatorname{ALG})]$, we can obtain a high-probability result for maximizing monotone submodular functions under noise. The idea is to repeat Algorithm 1 multiple times and output the best solution. The challenge here is to compare two sets $S_{1},S_{2}$ using noisy values, without access to the true values $f(S_{1}),f(S_{2})$. For monotone functions, one can construct another surrogate function $\tilde{f}_{0}(S)$, by randomly removing an element from $S$, to do the comparison. Monotonicity combined with submodularity ensures $f(S)\geq f_{0}(S)\geq(1-\frac{1}{|S|})f(S)$. Such an idea was also utilized by [16]. However, constructing such surrogate functions for non-monotone functions becomes technically challenging. The surrogate function above no longer works since removing an element could actually improve the objective. Indeed, an approximate version of monotonicity is not true in general. We leave for future work to obtain high-probability results for maximizing non-monotone functions under noise. See Appendix F for a detailed discussion.

### v2

@@ Abstract
We consider the problem of maximizing a submodular function with access to a noisy value oracle for the function instead of an exact value oracle. Similar to prior work [huang_efficient_2022, hassidim2017submodular], we assume that the noisy oracle is persistent in that multiple calls to the oracle for a specific set always return the same value. In this model, hassidim2017submodular design a (11/e)(1-1/e)-approximation algorithm for monotone submodular maximization subject to a cardinality constraint and huang_efficient_2022 design a (11/e)/2(1-1/e)/2-approximation algorithm for monotone submodular maximization subject to any arbitrary matroid constraint. In this paper, we design a meta-algorithm that allows us to take any “robust” algorithm for exact submodular maximization as a black box and transform it into an algorithm for the noisy setting while retaining the approximation guarantee. By using the meta-algorithm with the measured continuous greedy algorithm, we obtain a (11/e)(1-1/e)-approximation (resp. $1/e$-approximation) for monotone (resp. non-monotone) submodular maximization subject to a matroid constraint under noise. Furthermore, by using the meta-algorithm with the double greedy algorithm, we obtain a $1/2$-approximation for unconstrained (non-monotone) submodular maximization under noise.
@@ 1 Introduction
The standard model in submodular maximization assumes that we have value oracle access to the submodular function $f$ where, given a set $S$, we can retrieve the exact value of $f(S)$. However, in many settings it is not realistic to assume the existence of an exact value oracle. For example, in machine learning, we can never evaluate the true loss function of a model as we do not have access to the true distribution of the population. At best, we may have access to a noisy version of the loss function. In this paper, we study a model for noisy submodular maximization introduced by hassidim2017submodular where querying a set $S$ returns only an unbiased estimate of $f(S)$. A notable feature of this model is persistent noise, where querying the value of a set twice returns the same value. Thus, simply querying a submodular function multiple times at the same input cannot be used to denoise the function. Nonetheless, one can still ask whether existing algorithms work in this setting. However, hassidim2017submodular show that this is probably unlikely. For example, they show that the natural greedy algorithm is too sensitive to noise and only obtains an $o(1)$-approximation despite being the optimal algorithm when there is no noise. Thus, new algorithms are needed to deal with noise.
For the problem of maximizing a noisy monotone submodular function subject to a cardinality constraint, hassidim2017submodular designed an algorithm that achieves a tight (11/e)(1-1/e)-approximation. At a high-level, their algorithm computes a smooth surrogate function which essentially averages a few queries together to obtain an estimate that may have less noise but potentially be biased. They then apply the greedy algorithm on this surrogate function to obtain a (11/e)(1-1/e)-approximation. In a follow-up work, huang_efficient_2022 proved a (11/e)/2(1-1/e)/2-approximation ratio for noisy monotone submodular maximization under an arbitrary matroid. This approximation ratio can be improved to 11/e1-1/e if the matroid satisfies a “strongly base-orderable” property. While huang_efficient_2022 also make use of the smoothing technique, their algorithm is based on a local search algorithm [filmus2014monotone].
@@
While prior work has established that submodular maximization is feasible in the noisy setting, a notable downside is that the algorithms and analyses are designed for their specific problems. For example, hassidim2017submodular only obtain results for monotone submodular maximization subject to a cardinality constraint and, while huang_efficient_2022 extend to arbitrary matroids, their analyses have a factor of $2$ gap with the optimal result, despite an algorithm which is provably optimal in the non-noisy setting. Our main question is whether such specific analyses are necessary. In particular, we want to answer the following question:
@@
Our main contribution in this paper is an affirmative answer to the above question. In particular, we show that any algorithm for exact submodular optimization which is sufficiently “robust” can be translated to an algorithm for noisy submodular optimization with only an $o(1)$ loss in the approximation ratio. For many settings, this obviates the need for designing new algorithms and instead relies only on checking whether existing algorithms are sufficiently robust. Our technique builds on but differs from [hassidim2017submodular]: we propose the use of a random surrogate function, which circumvents the limitation of a deterministic surrogate function in [hassidim2017submodular].
For our first application, we instantiate our framework with the continuous greedy algorithm [calinescu_maximizing_2011, feldman_unified_2011]. Specifically, we use the measured continuous greedy algorithm of feldman_unified_2011 which gives a (11/e)(1-1/e)-approximation for monotone submodular maximization and a $1/e$-approximation for non-monotone submodular maximization, both subject to matroid constraints. As our framework inherits the approximation guarantees of these algorithms, we achieve the same approximation ratios for the noisy setting, with only $o(1)$ loss in the approximation ratio. Our second application is to instantiate our framework with the double greedy algorithm [Buchbinder_Feldman_Seffi_Schwartz_2015] to obtain a $1/2$-approximation for noisy unconstrained submodular maximization, which is tight even without noise. To our knowledge, we provide the first set of results for noisy submodular maximization for non-monotone functions, in both unconstrained and matroid constrained settings. We summarize our results in Table 1.
@@
Another work that is closely related to ours is hassidim2018optimization, which also studies the setting where the noise is persistent. In particular, they considered intersection of $P$ matroids and show a $1/(1+P)$-approximation. There have also been works that look at the noisy submodular setting but without persistent noise: for example, singla2016noisy consider a model where one only has preference information, and chen2023threshold assume that one has access to noisy marginal estimates instead of a value oracle.
Another related line of work is optimizing approximately submodular functions, where the noise is not unbiased and can be adversarial. horel2016maximization show that even there is only $1/\sqrt{n}$ noise, the problem becomes intractable and prove exponential lower bounds on the query complexity. On the other hand, they show that constant approximation algorithms are achievable when the “curvature” is bounded or when the noise is small in terms of the rank of the matroid. zheng2021maximizing study maximizing approximately $k$-submodular functions. chierichetti2022additive study functions that satisfy submodularity $\varepsilon$-approximately and convert them to exactly submodular functions that are $\varepsilon n^{2}$-close to the original functions, establishing a reduction to the exact submodular setting. A difference with our work is that our noisy value function is far from being $\varepsilon$-approximately submodular. A key contribution of our work is to convert our noisy value function to a surrogate function that is close to the underlying submodular function, which then enables a reduction similar to [chierichetti2022additive].
A less related line of work is on online submodular maximization [streeter2008online, roughgarden2018optimal, soma2019no, harvey2020improved, niazadeh2021online]. Here, one is given a sequence of submodular functions $f_{1},f_{2},\ldots$ and before seeing $f_{t}$, a player must decide on a set to play. The goal is to minimize $\alpha$-regret” which is the difference between the value obtained by the player and an $\alpha$-approximation of the optimum. Although the problem setting is different, these works also tend to show that many known algorithms are “robust” and develop a way to reuse these algorithms in the online setting. Such an approach is similar to ours as we show that a number of algorithms are robust and reuse them for the noisy setting.
@@ 2 Preliminaries
A noisy value oracle for $f$ is denoted by $\tilde{f}:2^{N}\to{\mathbb{R}}_{\geq 0}$. Following previous works [hassidim2017submodular, huang_efficient_2022], we consider a multiplicative noisy value oracle defined by $\tilde{f}(S)=\xi_{S}f(S)$ for all sets $S\subseteq N$, where $\xi_{S}$ is a non-negative random variable distributed according to some distribution $\mathcal{D}$. We call $\xi_{S}$ the noise multiplier for set $S$. Following [hassidim2017submodular, huang_efficient_2022], we assume that the noisy value oracle satisfies the following three properties:
@@
Definition 2.1 (see, e.g., wainwright_high-dimensional_2019).
@@
Sub-exponential distributions are a large class of distributions, including bounded, Gaussian, and exponential distributions. For example, a random variable $\xi$ bounded in $[0,B]$ is sub-exponential with parameters $(\nu=B,\alpha=0)$. Moreover, sub-exponential distributions include the generalized exponential tail distributions that have been considered by previous work on monotone submodular maximization under noise [hassidim2017submodular, huang_efficient_2022]. We assume that the parameters $(\nu,\alpha)$ of the sub-exponential distribution $\mathcal{D}$ are known while the distribution $\mathcal{D}$ itself is unknown.
@@ 3 A Unified Approach to Noisy Submodular Maximization
Our reduction uses an idea proposed by previous work on noisy submodular maximization [hassidim2017submodular, huang_efficient_2022]. This idea is to estimate the value $f(S)$ of a set $S$ using the noisy value of some surrogate function $F(S)=\frac{1}{|\mathcal{T}_{S}|}\sum_{T\in\mathcal{T}_{S}}f(T)$ where $\mathcal{T}_{S}$ is a collection of sets related to $S$. One example is to choose a small set $H$ and let $\mathcal{T}_{S}=\{S\cup H^{\prime}\text{ for $H^{\prime}\subseteq H$}\}$. In previous work, different surrogate functions were constructed for different settings. Here, we construct a single surrogate function for all the settings we consider (unconstrained non-monotone maximization, monotone and non-monotone maximization under matroid constraints); see Section 3.1 for details. Then, the reduction is to run the algorithm $\mathcal{A}$ to maximize the noisy surrogate function $\tilde{F}(S)=\frac{1}{|\mathcal{T}_{S}|}\sum_{T\in\mathcal{T}_{S}}\tilde{f}(T)\approx F(S)$.
The main technical challenge here is to ensure that the surrogate function $F(S)$ approximates the true function $f(S)$ well. As shown by [hassidim2017submodular, huang_efficient_2022], a deterministic surrogate function does not always guarantee a good approximation to the true function. So, they need to use arguments that are specific to their algorithms to show that the surrogate functions work well. A key insight in our work is the use of a random surrogate function. We show that such a random surrogate function can approximate the true function in all settings with matroid constraints (see Section 3.3).
@@ 3.1 Surrogate Function
Let $H\subseteq N$ be a subset of size $|H|=h$ where $h$ is a small integer. We call $H$ a smoothing set. Let $t<h$ be another integer. Let $H[t]=\{H^{\prime}\subseteq H:|H^{\prime}|=t\}$ be all the subsets of $H$ of size $t$; there are $\binom{h}{t}$ such subsets. We use $H^{\prime}\sim H[t]$ to denote sampling a subset $H^{\prime}\subseteq H$ of size $t$ uniformly at random. Define a surrogate function $F^{H,t}$ as follows:
@@ 3.2 A Meta-Algorithm for Noisy Submodular Maximization
Lemma 3.6 (donald_generalised_1991).
@@ 4.1 Matroid Constraints
We first consider maximizing monotone and non-monotone submodular functions under matroid constratins. We prove that the measured continuous greedy algorithm of feldman_unified_2011 is robust. Here, we recall the algorithm (with exact value oracle). First, define the multilinear extension of a submodular function $f$ as
@@
where the $\operatorname*{arg\,max}$ is taken over the matroid polytope $\mathcal{P}$. Then we update $x_{i}(t+\delta)=x_{i}(t)+\delta(1-x_{i}(t))y_{i}^{*}(t)$. Finally, we use pipage rounding [calinescu_maximizing_2011] (which is an obvlious rounding scheme) to convert the fractional solution $x(1)$ to a discrete set $S\in\mathcal{I}$. The pipage rounding technique guarantees that $\mathbb{E}[f(S)]\geq F(x(1))$. In addition, we have that $F(x(1))/f(\operatorname{OPT})\geq 1-1/e-O(n^{3}\delta)$ when $f$ is monotone and $F(x(1))/f(\operatorname{OPT})\geq 1/e-O(n^{3}\delta)$ when $f$ is non-monotone.
@@
For submodular function maximization under matroid constraint, the measured continuous greedy algorithm [feldman_unified_2011] obtains a (11/e)(1-1/e)-approximation when $f$ is monotone and a $(1/e)$-approximation when $f$ is non-monotone. Moreover, the algorithm is $O(n\varepsilon)$-robust.
@@
We apply Theorem 3.4 to the problem of maximizing monotone submodular functions with matroid constraints under a noisy value oracle. Let the algorithm $\mathcal{A}$ in Algorithm 1 be the measured continuous greedy algorithm [feldman_unified_2011] mentioned above. Fix parameter $\varepsilon\in(0,\frac{2\nu^{2}}{\alpha})$. Choose integer $m\geq\max\{2,8\nu^{2}\}\frac{n^{4}}{\varepsilon^{2}}\big(n+\log(4n)\big)=\tilde{O}(\frac{n^{5}}{\varepsilon^{2}})$, $t\geq\log_{2}(4m)=\Theta(\log(\frac{n}{\varepsilon}))$, and $h=t^{2}=\Theta(\log^{2}(\frac{n}{\varepsilon}))$. Then, we apply Theorem 3.4 and Lemma 4.1 with parameter $\varepsilon_{1}=\frac{\varepsilon f_{\max}}{n^{2}}$ to obtain:
@@
Fix $\varepsilon\in(0,\frac{2\nu^{2}}{\alpha})$. Suppose $n\geq\frac{1}{\varepsilon}$ and the matroid’s rank $r\geq\Omega(\frac{1}{\varepsilon}\log^{2}(\frac{n}{\varepsilon}))$. By letting the $\mathcal{A}$ in Algorithm 1 be the measured continuous greedy algorithm [feldman_unified_2011], we obtain a polynomial-time algorithm for maximizing monotone submodular functions under matroid constraints with noisy oracle satisfying $\mathbb{E}[f(\operatorname{ALG})]~\geq~\big(1-\frac{1}{e}-O(\varepsilon)\big)f(O^{*})$.
Previous work [huang_efficient_2022] gave a polynomial-time algorithm with $(1-1/e)/2-O(\varepsilon)$ approximation ratio for noisy monotone submodular maximization under matroid constraints, assuming $n\geq\varepsilon^{-4}$ and $r\geq\varepsilon^{-4/3}$. Our Theorem 4.2 improves [huang_efficient_2022] by increasing the approximation ratio to $1-1/e-O(\varepsilon)$ (which is the tight ratio even with the exact value oracle) as well as relaxing the condition on $n$ and $r$.
In the special case of cardinality constraints, [huang_efficient_2022]’s algorithm achieves $(1-1/e-O(\varepsilon))$ approximation with $\tilde{O}(r^{2}n^{3}/\varepsilon)$ query complexity (their Theorem 4.6). By letting $\mathcal{A}$ be the greedy algorithm with query complexity $Q(\mathcal{A})=O(rn)$, our algorithm achieves the same approximation ratio with query complexity $m\cdot Q(\mathcal{A})=\tilde{O}(n^{5}/\varepsilon^{2})\cdot O(rn)=\tilde{O}(rn^{6}/\varepsilon^{2})$. While our query complexity might be worse than [huang_efficient_2022], the advantage of our algorithm lies in its generality and better approximation ratio in the more general matroid constraints case.
@@
Fix $\varepsilon\in(0,\frac{2\nu^{2}}{\alpha})$. Suppose $n\geq\frac{1}{\varepsilon}$ and the matroid’s rank $r\geq\Omega(\frac{1}{\varepsilon}\log^{2}(\frac{n}{\varepsilon}))$. By letting the $\mathcal{A}$ in Algorithm 1 be the measured continuous greedy algorithm [feldman_unified_2011], we obtain a polynomial-time algorithm for maximizing non-monotone submodular functions under matroid constraints with noisy oracle satisfying $\mathbb{E}[f(\operatorname{ALG})]~\geq~\big(\frac{1}{e}-\frac{1}{\Theta(\log\frac{n}{\varepsilon})}-O(\varepsilon)\big)f(O^{*})$.
The $\frac{1}{e}\approx 0.367$ approximation above is not necessarily tight. buchbinder2024constrained design a $0.401$-approximation algorithm for non-monotone submodular maximization under matroid constraint with exact value oracle. Their algorithm is quite technical, so we leave as an open question whether their algorithm is robust against approximate value oracle. If their algorithm is robust, then it can be directly converted to an algorithm for noisy non-monotone submodular maximization under matroid constraint achieving 0.401o(1)0.401-o(1) approximation, using our Algorithm 1.
@@
While our main results (Theorems 3.4, 4.2 and 4.3) are stated in terms of the expected value $\mathbb{E}[f(\operatorname{ALG})]$, we can obtain a high-probability result for maximizing monotone submodular functions under noise. The idea is to repeat Algorithm 1 multiple times and output the best solution. The challenge here is to compare two sets $S_{1},S_{2}$ using noisy values, without access to the true values $f(S_{1}),f(S_{2})$. For monotone functions, one can construct another surrogate function $\tilde{f}_{0}(S)$, by randomly removing an element from $S$, to do the comparison. Monotonicity combined with submodularity ensures $f(S)\geq f_{0}(S)\geq(1-\frac{1}{|S|})f(S)$. Such an idea was also utilized by [huang_efficient_2022]. However, constructing such surrogate functions for non-monotone functions becomes technically challenging. The surrogate function above no longer works since removing an element could actually improve the objective. Indeed, an approximate version of monotonicity is not true in general. We leave for future work to obtain high-probability results for maximizing non-monotone functions under noise. See Appendix F for a detailed discussion.